# Corks, exotic 4-manifolds and genus functions

### v3

@@ Abstract.
We prove that every 4-dimensional oriented handlebody without 3- and 4-handles can be modified to admit infinitely many exotic smooth structures, and further that their genus functions are pairwise equivalent. We furthermore show that, for any 4-manifold admitting an embedding into a symplectic 4-manifold with weakly convex boundary, its genus function is algebraically realized as those of infinitely many pairwise exotic 4-manifolds. We also prove that, for any two 4-manifolds with non-degenerate intersection forms, algebraic inequivalences of genus functions are stable under connected sums and boundary sums with a certain type of 4-manifolds having arbitrarily large second Betti numbers. Furthermore, we introduce a notion of genus function type for diffeomorphism invariants and show that similar results hold for any such invariant. As an application of our exotic 4-manifolds, we also prove that for any (possibly non-orientable) 4-manifold, every submanifold of codimension less than two satisfying a mild condition can be modified to admit infinitely many exotically knotted copies.
@@
4-manifolds; genus functions; corks; knotted submanifolds; cobordisms
@@ 1. Introduction
Lawson conjectured in his 1997 survey [49] that, for simply connected compact 4-manifolds, genus functions determine smooth structures. Indeed, for rational surfaces and ruled surfaces, genus functions uniquely determine their smooth structures among symplectic 4-manifolds (see Proposition 2.6). It later turned out that his conjecture is false ([8], [36]), but the counterexamples are not interesting, because their second homology groups are trivial, and hence the underlying topological structures uniquely determine genus functions. Thus, the conjecture remains intriguing, especially for 4-manifolds with large second Betti numbers. It would also be natural to ask whether a 4-manifold can simultaneously admit exotic smooth structures with the same genus functions and those with distinct genus functions.
Let us recall that the genus function of a connected oriented smooth 4-manifold $X$ is defined to be the map $H_{2}(X)\to\mathbb{Z}$ which sends a second homology class to the minimal genus of a smoothly embedded surface representing the class. Genus functions of two 4-manifolds will be called equivalent if there exists a homeomorphism between them whose induced isomorphism preserves the genus functions. They will be called algebraically equivalent if there exists an isomorphism between their second homology groups that preserves the genus functions and the intersection forms.
In this paper, we prove that every 4-manifold with boundary satisfying mild conditions can be modified to admit infinitely many exotic smooth structures. We then show that exotic 4-manifolds with equivalent genus functions exist in abundance. Furthermore, we study behavior of genus functions and properties of related invariants. We also produce exotically knotted submanifolds.
@@
Throughout this paper, all manifolds, submanifolds, boundaries, handlebodies and surfaces are assumed to be compact, connected, oriented and smooth unless otherwise stated. Also, two 4-manifolds are said to be exotic to each other if one is orientedly homeomorphic to but not unorientedly diffeomorphic to the other.
@@ 1.1. Exotic 4-manifolds and genus functions
To state our results shortly, we introduce the following terminology.
@@
For two 4-manifolds $X,Y$ having non-empty (possibly disconnected) boundary, we will say that $Y$ topologically resembles $X$ if $Y$ is homotopy equivalent to $X$, the intersection form of $Y$ is isomorphic to that of $X$, and $\partial Y$ is homology cobordant to $\partial X$.
Let us recall that a $2$-handlebody is a handlebody consisting of finitely many handles with indices not greater than two. They form a very large class of 4-manifolds with boundary, realizing arbitrary large integers as their second Betti numbers $b_{2}$. (Beware that in some papers, 2-handlebodies mean handlebodies consisting of only 0- and 2-handles, which form a strictly smaller class than ours.) We prove that every 4-dimensional 2-handlebody can be modified to admit infinitely many exotic smooth structures with pairwise equivalent genus functions, by utilizing corks.
@@
For any $4$-dimensional $2$-handlebody $X$, there exist infinitely many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise equivalent.
Even if the condition on genus functions is removed, this result is still new and shows the existence of infinitely many exotic smooth structures for a very large class of 4-manifolds, strongly improving the following earlier results: every 2-handlebody with non-trivial second homology group can be modified to admit arbitrarily many exotic smooth structures (Akbulut and the author [11]); there exist contractible 4-manifolds admitting infinitely many exotic smooth structures (Gompf [36]).
@@
We furthermore prove that various genus functions can not determine smooth structures.
@@
Suppose that a $4$-manifold $X$ admits an embedding into a symplectic $4$-manifold with weakly convex boundary. Then, there exist infinitely many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise equivalent and are algebraically equivalent to that of $X$.
The assumption of this theorem holds for a large class of smooth 4-manifolds. For example, blow ups of 4-manifolds admitting Stein structures and their codimension-00 submanifolds satisfy the assumption. Furthermore, these 4-manifolds provide varieties of genus functions even within fixed homeomorphism types ([11], [68], [70]). We here state only simple examples.
@@
$(1)$ For each positive integer $n$, there exist infinitely many pairwise exotic simply connected $4$-manifolds ((((with boundary)))) such that their genus functions are pairwise equivalent and are algebraically equivalent to that of $n\mathbb{C}P^{2}$.
$(2)$ If $X$ is the total space of a $D^{2}$-bundle over a closed surface, then there exist infinitely many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise equivalent and are algebraically equivalent to that of $X$.
@@
There exist infinitely many pairwise exotic simply connected $4$-manifolds with pairwise algebraically inequivalent genus functions such that each of their genus functions is equivalent to those of another infinitely many pairwise exotic $4$-manifolds.
@@
For any $4$-dimensional $2$-handlebody $X$ with non-trivial second homology group, there exist arbitrarily many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise algebraically inequivalent for any orientations and that each of their genus functions is equivalent to those of infinitely many pairwise exotic $4$-manifolds.
Exotic 4-manifolds in Theorems 1.3, 1.5 and 1.8 have more properties inherited from a given 4-manifold $X$, see Theorems 9.4, 9.5 and 9.8 for details.
It remains interesting to ask whether exotic closed 4-manifolds with equivalent genus functions exist. For closed 4-manifolds, we instead show that algebraic equivalent classes do not determine homeomorphism types even within 4-manifolds having isomorphic cohomology rings and that there are many such 4-manifolds, realizing arbitrarily large $b_{2}$. See Proposition 8.13 for details. During the preparation of this paper, Stipsicz and Szabó [61] proved that infinitely many pairwise non-diffeomorphic closed 4-manifolds with the cohomology ring of $S^{2}\times S^{2}$ obtained in [29] have pairwise algebraically equivalent genus functions. It is not known whether their 4-manifolds are pairwise homeomorphic. They showed algebraic equivalence by explicitly determining all the values of the genus functions, but our proof is completely different from their approach.
After the first version of this paper appeared on arXiv on January 2025, Nakamura submitted the paper [53] to arXiv which gives a different example of Corollary 1.4.(1) in the case where the intersection form is represented by the $1\times 1$ matrix 00.
@@ 1.2. Stabilities of algebraic inequivalences of genus functions
To show that our exotic 4-manifolds have equivalent genus functions, we prove that attachment of a cobordism of certain type preserves genus functions (See Sections 6 and 7). As an application, we discuss behavior of genus functions under connected sums and boundary sums with a certain type of 4-manifolds.
For an $S^{2}$-link $L$ in $S^{4}$, let $S^{4}(L)$ denote the closed 4-manifold obtained from $S^{4}$ by performing the surgery on $L$, that is, by removing the interior of the tubular neighborhood $S^{2}\times D^{2}$ for each component of $L$ and gluing $S^{1}\times D^{3}$ along the resulting boundary. A simple example of $S^{4}(L)$ is $n(S^{1}\times S^{3})$. Each $S^{4}(L)$ has varieties of codimension-00 submanifolds having arbitrary large $b_{2}$ (see Section 8). For example, $\Sigma_{g}\times D^{2}$ $(g\geq 0)$ with 2-handles attached along a strongly slice link with 0-framings are such examples, where $\Sigma_{g}$ denotes the closed surface of genus $g$.
In the rest of this subsection, for each $i=1,2$, let $X_{i}$ be a ((((possibly closed)))) 4-manifold, $L_{i}$ be an $S^{2}$-link in $S^{4}$, and $Z_{i}$ be a connected summand or a codimension-00 submanifold of $S^{4}(L_{i})$.
We will prove that the values of the genus functions of 4-manifolds are preserved under connected sums and boundary sums with $Z_{i}$’s, see Theorem 8.4. Applying this result, we show that algebraic inequivalences of genus functions are stable under these operations.
@@
This result contrasts with the fact that the Seiberg–Witten invariants of closed 4-manifolds vanish after taking a connected sum with $S^{1}\times S^{3}$ and hence cannot detect them after the connected sum.
Even when $Z_{i}$’s have non-trivial $b_{2}$, a similar stability still holds under additional assumptions.
@@
We also give similar results without the above torsion-free assumption, see Theorems 8.10 and 8.11. These stabilities for algebraic inequivalences show that for any family of exotic 4-manifolds with algebraically inequivalent genus functions, the family remain (orientedly) exotic under connected sums and boundary sums with homeomorphic 4-manifolds satisfying the assumption. We can apply these stabilities to various families of exotic 4-manifolds having algebraically inequivalent genus functions (e.g. [11], [67], [68], [70]). Furthermore, we give similar stabilities for attachments of certain cobordisms, see Sections 6 and 7.
@@
Note that the summands $Y_{i}$’s are not necessarily pairwise diffeomorphic. As seen from the proof, we can realize various homeomorphism types as those of $X_{i}$’s. Although the above theorem assumes that each $Y_{i}$ admits a Stein structure, the above theorem holds under the weaker condition that each $Y_{i}$ admits an embedding into a symplectic 4-manifold with weakly convex boundary. This will be discussed in the paper [72] announced in [71]. Moreover, in [72], we will produce various families of exotic 4-manifolds which are stably exotic under connected sums with arbitrary closed definite 4-manifolds.
@@ 1.3. Diffeomorphism invariants of genus function type
Furthermore, we show that an invariant of genus function type yields lower bounds for values of genus functions, which can be sharp depending on the invariant, see Proposition 11.8. We also prove stabilities of algebraic inequivalences similar to Theorems 1.9 and 1.10, see Theorems 11.12 and 11.13. These stabilities imply that if (algebraic equivalent classes of) an invariant of genus function type can detect a family of exotic 4-manifolds, then the invariant can detect infinitely many families of exotic 4-manifolds, see Corollary 11.14. By contrast, we also prove that there exist various families of exotic 4-manifolds which cannot be detected by any invariant of genus function type, see Theorem 11.15. We will give further related results in Section 11.
@@ 1.4. Exotically knotted submanifolds
Applying our exotic 4-manifolds, we also discuss exotically knotted submanifolds. It is natural to ask whether a given submanifold of a 4-manifold admits exotically knotted copies and/or exotically embedded copies, and many examples of submanifolds admitting such copies have been constructed, see, for example, [26], [27], [11], [14], [59], [40], [52], [45] and the references therein. Here a pair of submanifolds of a manifold will be called exotically knotted if they are diffeomorphic, topologically ambient isotopic, but not smoothly ambient isotopic. Also, a pair of submanifolds will be called exotically embedded if they are diffeomorphic, the ambient manifold admits a self-homeomorphism that maps one to the other, but the ambient manifold does not admit such a self-diffeomorphism. We furthermore require natural conditions for orientations and boundaries as well (see Definition 10.1), but our results still hold under the above simpler conditions.
@@
For a codimension-00 submanifold of a 4-manifold, the closure of the complement will be called the exterior of the submanifold. Utilizing our exotic 4-manifolds, we show that every codimension-00 submanifold satisfying a mild condition can be modified to admit infinitely many exotically knotted and exotically embedded copies.
@@
Let $X$ be a codimension-00 submanifold of a $4$-manifold $Z$. These manifolds are possibly non-orientable and may have ((((possibly disconnected)))) boundaries. Suppose the exterior of $X$ is either a ((((orientable)))) $2$-handlebody or diffeomorphic to a submanifold of a symplectic $4$-manifold with weakly convex boundary. Then $Z$ admits infinitely many pairwise exotically knotted and exotically embedded codimension-00 submanifolds topologically resembling $X$.
Exotically knotted submanifolds in this theorem have more properties inherited from a given submanifold $X$, see Theorem 10.2 for details. This result improves an earlier result of Akbulut and the author [11], which modified $X$ into arbitrarily many exotically embedded submanifolds additionally assuming that the exterior is a 2-handlebody with non-trivial second homology group. Note that their examples were not shown to be exotically knotted in general.
@@
Every ((((possibly non-orientable)))) $4$-manifold admits infinitely many pairwise exotically knotted and exotically embedded ((((possibly non-orientable)))) submanifolds of codimension-00.
We also discuss exotically knotted codimension-1 submanifolds by applying Theorem 1.12. Prior results in the literature are realizations of many 3-manifolds as exotically knotted codimension-1 submanifolds of some orientable 4-manifolds (see e.g. [40] and the references therein). By contrast, we show that for a given (possibly non-orientable) 4-manifold, every codimension-$1$ submanifold satisfying a mild condition can be modified to admit infinitely many exotically knotted (and exotically embedded) copies, using homology cobordisms embedded in the given 4-manifold.
@@
Let $Z$ be a ((((possibly non-orientable)))) $4$-manifold, and let $M$ be a closed $3$-manifold embedded in the interior of $Z$ as a separating submanifold. Suppose the exterior of the tubular neighborhood of $M$ has a connected component which is either a ((((orientable)))) $2$-handlebody or diffeomorphic to a submanifold of a symplectic $4$-manifold with weakly convex boundary. Then $Z$ admits infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds which are homology cobordant to $M$.
As corollaries, we show that there exist infinitely many closed $3$-manifolds such that every ((((possibly non-orientable)))) $4$-manifold realizes each of them as infinitely many exotically knotted submanifolds.
@@
For any closed $3$-manifold embedded in $S^{4}$, every ((((possibly non-orientable)))) $4$-manifold admits infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds which are homology cobordant to the $3$-manifold.
@@
There exist infinitely many pairwise not homology cobordant closed $3$-manifolds such that, for each $3$-manifold of them, every ((((possibly non-orientable)))) $4$-manifold admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to the $3$-manifold.
Note that many closed 3-manifolds do not have the above property, since there are many closed 3-manifolds not admitting any embedding into $S^{4}$ (see e.g. [41] and the references therein).
For 1-manifolds embedded in a 4-manifold, the existence of a topological ambient isotopy guarantees that of a smooth ambient isotopy, so there exist no exotically knotted 1-manifolds. It would be natural to ask whether the above corollary holds for other codimensions.
Question 1.17.

### v5

@@ Abstract.
We prove that every 4-dimensional oriented handlebody without 3- and 4-handles can be modified to admit infinitely many exotic smooth structures, and moreover prove that their genus functions are pairwise equivalent. We furthermore show that for any 4-manifold admitting an embedding into a symplectic 4-manifold with weakly convex boundary, its genus function is algebraically realized as those of infinitely many pairwise exotic 4-manifolds. In addition, we prove that algebraic inequivalences of genus functions are stable under connected sums and boundary sums with a certain type of 4-manifolds having arbitrarily large second Betti numbers. Besides, we introduce a notion of genus function type for diffeomorphism invariants, and show that any such invariant shares properties similar to all the preceding results and yields lower bounds for the values of genus functions. As an application of our exotic 4-manifolds, we also prove that for any (possibly non-orientable) 4-manifold, every submanifold of codimension at most one satisfying a mild condition can be modified to admit infinitely many exotically knotted copies.
@@
4-manifolds; genus functions; corks; knotted submanifolds
@@ 1. Introduction
Lawson conjectured in his 1997 survey [51] that, for simply connected compact 4-manifolds, genus functions determine smooth structures. Indeed, for rational surfaces and ruled surfaces, genus functions uniquely determine their smooth structures among symplectic 4-manifolds (Proposition 2.7). It later turned out that his conjecture is false ([8], [37]), but the counterexamples are not particularly interesting, because their second homology groups are trivial, and hence the underlying topological structures uniquely determine genus functions. Thus, the conjecture remains intriguing especially for 4-manifolds with large second Betti numbers, and naturally raises the problem whether a 4-manifold can simultaneously admit exotic structures with the same genus function and those with distinct genus functions.
Let us recall that the genus function of a connected oriented smooth 4-manifold $X$ is defined to be the map $g_{X}:H_{2}(X)\to\mathbb{Z}$ which sends a class $\alpha$ to the minimum genus of a smoothly embedded surface representing $\alpha$. Genus functions of two 4-manifolds will be called equivalent if there exists a homeomorphism between them whose induced isomorphism preserves the genus functions. They will be called algebraically equivalent if there exists an isomorphism between their second homology groups that preserves their genus functions and intersection forms.
In this paper, we prove that every 4-manifold with boundary satisfying mild conditions can be modified to admit infinitely many exotic smooth structures. Applying this result, we answer the aforementioned problems on genus functions and also modify submanifolds of 4-manifolds into exotically knotted ones. Furthermore, we reveal several interesting properties of genus functions. Besides, we introduce a notion of genus function type for diffeomorphism invariants of 4-manifolds, and show that any such invariant shares many properties of genus functions.
@@
Throughout this paper, unless otherwise stated, all manifolds, submanifolds, boundaries, handlebodies, and surfaces are assumed to be compact, connected, oriented, and smooth, and all embeddings of manifolds are assumed to be smooth but not necessarily proper. A pair of 4-manifolds is called exotic (resp. orientedly exotic) if they are orientedly homeomorphic but not unorientedly diffeomorphic (resp. not orientedly diffeomorphic). Here, “orientedly” and “unorientedly” respectively mean “orientation-preserving” and “not necessarily orientation-preserving”.
We require exotic 4-manifolds to be orientedly homeomorphic, since this assumption is necessary to ensure that their connected sums and boundary sums with a common 4-manifold always yield homeomorphic 4-manifolds. We note that boundary 3-manifolds of two exotic 4-manifolds are always orientedly diffeomorphic.
@@ 1.1. Exotic 4-manifolds and genus functions
For convenience, we introduce a notion of topological similarity for 4-manifolds, which we call HIHC-equivalence.
@@
Two 4-manifolds with (possibly disconnected) boundary are called HIHC-equivalent if they are homotopy equivalent, their intersection forms are isomorphic, and their boundaries are homology cobordant.
We note that two orientedly homeomorphic 4-manifolds are necessarily HIHC-equivalent, and the converse also holds for simply connected closed (smooth) 4-manifolds due to Freedman’s celebrated theorem.
Let us recall that a $2$-handlebody is a handlebody consisting of finitely many handles whose indices are at most two. $2$-handlebodies form a very large class of 4-manifolds with boundary, realizing arbitrarily large integers as their second Betti numbers $b_{2}$. (Note that in some papers, 2-handlebodies mean handlebodies consisting only of 0- and 2-handles, which form a much smaller class.)
We prove that every 4-dimensional 2-handlebody can be modified to admit infinitely many exotic smooth structures by using corks, and moreover show that these exotic structures have pairwise equivalent genus functions.
@@
For any $4$-dimensional 22-handlebody $X$, there exist infinitely many pairwise exotic $4$-manifolds which are HIHC-equivalent to $X$. Moreover, they can be chosen so that their genus functions are pairwise equivalent.
Even without the conclusion on genus functions, this result is new and shows the existence of infinitely many exotic smooth structures for a very large class of 4-manifolds, substantially strengthening the following earlier results: every 2-handlebody with non-trivial second homology group can be modified to admit arbitrarily many exotic structures (Akbulut and the author [11]); there exist contractible 4-manifolds admitting infinitely many exotic structures (Gompf [37]).
@@
We furthermore show that a given 4-manifold embedded in a symplectic $4$-manifold with weakly convex boundary can be modified to admit infinitely many exotic smooth structures. We moreover prove that the genus function of the given 4-manifold is algebraically realized as their genus functions, showing that there are various genus functions which cannot determine smooth structures.
@@
For any $4$-manifold $X$ admitting an embedding into a symplectic $4$-manifold with weakly convex boundary, there exist infinitely many pairwise exotic $4$-manifolds which are HIHC-equivalent to $X$. Moreover, they can be chosen so that their genus functions are pairwise equivalent and are algebraically equivalent to that of $X$.
The assumption of this theorem holds for a large class of smooth 4-manifolds. For example, blow ups of 4-manifolds admitting Stein structures and their codimension-0 submanifolds satisfy the assumption. Furthermore, these 4-manifolds provide varieties of genus functions even within fixed homeomorphism types ([11], [70], [72]). We here state only simple examples.
@@
$(1)$ For each positive integer $n$, there exist infinitely many pairwise exotic simply connected $4$-manifolds ((with boundary)) such that their genus functions are pairwise equivalent and are algebraically equivalent to that of $n\mathbb{C}P^{2}$.
$(2)$ For any $D^{2}$-bundle over a closed surface, there exist infinitely many pairwise exotic $4$-manifolds HIHC-equivalent to the total space $X$ of the bundle such that their genus functions are pairwise equivalent and are algebraically equivalent to that of $X$.
@@
There exist infinitely many pairwise exotic simply connected $4$-manifolds with pairwise algebraically inequivalent genus functions such that, for each $4$-manifold $X$ among them, there exist infinitely many pairwise exotic $4$-manifolds homeomorphic to $X$ whose genus functions are all equivalent to that of $X$.
@@
For any $4$-dimensional 22-handlebody $X$ with non-trivial second homology group, there exist arbitrarily many pairwise exotic $4$-manifolds HIHC-equivalent to $X$ and having pairwise algebraically inequivalent genus functions such that, for each $4$-manifold $X^{\prime}$ among them, there exist infinitely many pairwise exotic $4$-manifolds homeomorphic to $X^{\prime}$ whose genus functions are all equivalent to that of $X^{\prime}$.
We remark that exotic 4-manifolds in Theorems 1.3, 1.5 and 1.8 have more properties inherited from a given 4-manifold $X$, see Theorems 9.4, 9.5 and 9.8.
It remains interesting to see whether exotic closed 4-manifolds with equivalent genus functions exist. For closed 4-manifolds, we instead show that algebraic equivalent classes do not determine homeomorphism types even within 4-manifolds having isomorphic cohomology rings and that there are many such 4-manifolds, realizing arbitrarily large $b_{2}$. See Theorem 8.12 for details. During the preparation of this paper, Stipsicz and Szabó [63] proved that infinitely many pairwise non-diffeomorphic closed 4-manifolds with the cohomology ring of $S^{2}\times S^{2}$ obtained in [32] have pairwise algebraically equivalent genus functions. It is not known whether their 4-manifolds are pairwise homeomorphic. They showed algebraic equivalence by explicitly determining all the values of the genus functions, but our proof is completely different from their approach.
We remark that, after the first version of this paper appeared on arXiv on January 2025, Nakamura submitted the paper [56] to arXiv which gives an example of Corollary 1.4.(1) in the case where the intersection form is represented by the $1\times 1$ matrix 0, by using a different method.
@@ 1.2. Stabilities of algebraic inequivalences of genus functions
To show that our exotic 4-manifolds have equivalent genus functions, we prove that attachments of a class of cobordisms called quasi-invertible cobordisms always preserve the genus functions, which is of independent interest (see Sections 6 and 7). As an application, we discuss behavior of genus functions under connected sums and boundary sums with a certain type of 4-manifolds.
For an $S^{2}$-link $L$ in $S^{4}$, let $S^{4}(L)$ denote the closed 4-manifold obtained from $S^{4}$ by performing the surgery on $L$, that is, by removing the interior of the tubular neighborhood $S^{2}\times D^{2}$ for each component of $L$ and gluing $S^{1}\times D^{3}$ along the resulting boundary. A simple example of $S^{4}(L)$ is $n(S^{1}\times S^{3})$. As we will show in Subsection 8.1, each $S^{4}(L)$ has rich examples of codimension-0 submanifolds, realizing arbitrarily large $b_{2}$. For instance, $\Sigma_{g}\times D^{2}$ $(g\geq 0)$ with 2-handles attached along a strongly slice link (see Definition 6.6) with 0-framings are such examples, where $\Sigma_{g}$ denotes the closed surface of genus $g$.
In the rest of this subsection, for each $i=1,2$, let $X_{i}$ be a ((possibly closed)) 4-manifold, $L_{i}$ be an $S^{2}$-link in $S^{4}$, and $Z_{i}$ be a connected summand or a codimension-0 submanifold of $S^{4}(L_{i})$.
We will prove that the values of the genus functions of 4-manifolds are preserved under connected sums and boundary sums with $Z_{1}$ and $Z_{2}$ (Theorem 8.4). Applying this result, we show that algebraic inequivalences of genus functions are stable under these operations. In the case where $H_{2}(Z_{i})=0$, we prove the following theorem.
@@
This result contrasts with the fact that the Seiberg–Witten invariants of closed symplectic 4-manifolds vanish after taking a connected sum with $S^{1}\times S^{3}$ and hence cannot detect them after the connected sum.
In the general case, we prove similar stabilities under additional assumptions.
@@
Furthermore, we give similar results and refinements without assuming the torsion-free conditions (Theorems 8.9 and 8.10). However, the above theorem does not hold if the non-degeneracy assumption is removed (Remark 8.11).
These stabilities for algebraic inequivalences show that for any family of exotic 4-manifolds with algebraically inequivalent genus functions, the family remain (orientedly) exotic under connected sums and boundary sums with homeomorphic 4-manifolds satisfying the assumptions. One can apply these stabilities to various families of exotic 4-manifolds (e.g. [11], [69], [70], [72]) and produce further examples of exotic families by connected sums and boundary sums. Furthermore, we give similar stabilities for attachments of quasi-invertible cobordisms, see Subsection 7.2. We remark that the proofs of these stabilities use no gauge theory.
@@
In fact, we realize various homeomorphism types including not simply connected ones as those of $X_{1},\dots,X_{n}$, see Theorem 9.9. Note that the summands $Y_{1},\dots,Y_{n}$ can be taken as pairwise exotic 4-manifolds due to a result of [11] (see also Theorem 3.7). Although the above theorem assumes that each $Y_{i}$ admits a Stein structure, the above theorem holds under the weaker condition that each $Y_{i}$ admits an embedding into a symplectic 4-manifold with weakly convex boundary. This will be discussed in the paper [74] announced in [73]. Moreover, in [74], we will produce various families of exotic 4-manifolds which are stably exotic under connected sums with arbitrary closed definite 4-manifolds.
@@ 1.3. Diffeomorphism invariants of genus function type
Furthermore, we show that every invariant of genus function type yields lower bounds for values of genus functions, which can be sharp depending on the invariant, see Proposition 11.7. We also prove stabilities of algebraic inequivalences similar to Theorems 1.9 and 1.10, see Theorems 11.12 and 11.13. These stabilities imply that if (algebraic equivalent classes of) an invariant of genus function type can detect a family of orientedly exotic 4-manifolds, then the invariant can detect infinitely many families of orientedly exotic 4-manifolds, see Corollary 11.14. By contrast, we also prove that there exist various families of exotic 4-manifolds which can not be detected by any invariant of genus function type, see Theorems 11.15 and 11.16. We will give further related results in Section 11.
@@ 1.4. Exotically knotted submanifolds
Applying our exotic 4-manifolds, we also discuss exotically knotted and exotically embedded submanifolds. Here a pair of submanifolds of a manifold will be called exotically knotted if they are diffeomorphic, topologically ambient isotopic, but not smoothly ambient isotopic. Also, a pair of submanifolds will be called exotically embedded if they are diffeomorphic, the ambient manifold admits a self-homeomorphism that maps one to the other, but the ambient manifold does not admit such a self-diffeomorphism. We furthermore require natural conditions for orientations and boundaries as well (see Definition 10.1), but our results still hold under the above simpler conditions.
@@
It is a major problem in 4-dimensional topology to see whether a given submanifold of a 4-manifold admits exotically knotted and/or exotically embedded copies. Many examples of submanifolds admitting such copies have been constructed (see, for example, [27], [30], [11], [15], [61], [41], [55], [46] and the references therein), but this problem remains open in general. Utilizing our exotic 4-manifolds, here we show that every codimension-0 submanifold satisfying a mild condition on the exterior (i.e., the closure of the complement) can be modified to admit infinitely many exotically knotted and exotically embedded copies.
@@
Let $X$ be a $4$-manifold embedded in a $4$-manifold $Z$. These manifolds may be non-orientable and may have ((possibly disconnected)) boundaries. Suppose the exterior of $X$ either is a ((orientable)) 22-handlebody or admits an embedding into a symplectic $4$-manifold with weakly convex boundary, and assume the boundary of the exterior is connected. Then, $Z$ admits infinitely many pairwise exotically knotted and exotically embedded codimension-0 submanifolds which are HIHC-equivalent to $X$.
Furthermore, we show that each of these exotically knotted submanifolds contains the given submanifold $X$ as a submanifold and admits a deformation retraction onto $X$, indicating that the modification to $X$ is small, see Theorem 10.2 and also Remark 10.5. Note that the modification is necessary, since the 4-ball admits neither exotically knotted nor exotically embedded copies due to the disk theorem. The above theorem improves an earlier result of Akbulut and the author [11], which modified $X$ into arbitrarily many exotically embedded submanifolds additionally assuming that the exterior is a 2-handlebody with non-trivial second homology group. Their examples were not shown to be exotically knotted in general.
@@
Every ((possibly non-orientable)) 44-manifold admits infinitely many ((possibly non-orientable)) codimension-0 submanifolds which are pairwise exotically knotted and exotically embedded.
We also discuss codimension-1 submanifolds by applying Theorem 1.12. Prior results in the literature are realizations of many 3-manifolds as exotically knotted codimension-1 submanifolds of some orientable 4-manifolds (see e.g. [41] and the references therein). By contrast, we show that for a given (possibly non-orientable) 4-manifold, every codimension-$1$ separating submanifold satisfying a mild condition can be modified to admit infinitely many exotically knotted (and exotically embedded) copies, using homology cobordisms embedded in the given 4-manifold.
@@
Let $M$ be a closed $3$-manifold embedded in the interior of a ((possibly non-orientable)) 44-manifold $Z$ which may have ((possibly disconnected)) boundary. Suppose the exterior of the tubular neighborhood of $M$ has a connected component with connected boundary which either is a $4$-dimensional ((orientable)) 22-handlebody or admits an embedding into a symplectic $4$-manifold with weakly convex boundary. Then, $Z$ admits infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds which are homology cobordant to $M$.
As corollaries, we show that there exist infinitely many closed $3$-manifolds such that every ((possibly non-orientable)) 44-manifold realizes each of them as infinitely many exotically knotted submanifolds.
@@
For any closed $3$-manifold $M$ embedded in $S^{4}$, every ((possibly non-orientable)) 44-manifold admits infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds which are homology cobordant to $M$.
@@
There exist infinitely many pairwise not homology cobordant closed $3$-manifolds such that, for each $3$-manifold $M$ among them, every ((possibly non-orientable)) 44-manifold admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to $M$.
Note that many closed 3-manifolds do not have the above property. Indeed, there are many closed 3-manifolds not admitting any embedding into $S^{4}$ (see e.g. [42] and the references therein).
For a given closed 3-manifold $M$, we consider the question of what is the minimum second Betti number of a simply connected closed 4-manifold which contains a pair of exotically knotted (or exotically embedded) submanifolds diffeomorphic to $M$ (cf. [2, Definition 2.2], [41, Question 7.7]). This question naturally arises from the well-known fact that every closed 3-manifold admits an embedding into a simply connected closed 4-manifold (see e.g. [2, Theorem 2.1]). Here we show that every non-negative even integer is realized as such a minimum second Betti number by using Theorem 1.14.
Corollary 1.17.
For each even integer $n\geq 0$, there exist infinitely many pairwise not homology cobordant closed $3$-manifolds such that, for each $3$-manifold $M$ among them, a simply connected closed $4$-manifold with $b_{2}=n$ contains infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to $M$, but no simply connected closed $4$-manifold with $b_{2}<n$ contains a submanifold diffeomorphic to $M$.
In fact, we show that the above $4$-manifold with $b_{2}=n$ can be taken as $\frac{n}{2}(S^{2}\times S^{2})$, independently of the choice of the 3-manifold $M$.
We close this subsection by posing a question. As is well-known, every 4-manifold admits no pair of exotically knotted codimension-3 submanifolds. By contrast, Corollary 1.16 gives closed 3-manifolds such that every 4-manifold admits pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to them. Thus, the following question naturally arises for the remaining codimensions.
Question 1.18.